rendell father is 9times as old as he is.one year before he was 2years more that eleven times as old as rendell.find rendell present age.
Answer by mananth(16949)

let rendell's age be x
father's age = 9x.
..
1 year ago their ages were
x-1
9x-1
..
9x-1=11(x-1)+2
9x-1=11x-11+2
9x-11x=-11+1+2
-2x=-8
x=4 Rendell's age now
father's age = 9*4=36 years